If

G

â

ƒ

=

ã

5

t

ã

2

i

+

t

j

t

3

k

G⃗=〖5t〗2i+tj−t3k and

F

â

ƒ

=

s

i

n

t

i

â

c

o

s

t

j

F⃗=sinti–costj, what is

d

/

d

t

(

G

×

F

â

ƒ

)

d/dt(G×F⃗)


1
Expert's answer
2021-05-20T11:41:57-0400

"G=-2i+tj-3tk"

"F=isint+jcost"

"G\times F=\begin{vmatrix}
 i & j&k \\
 -2 & t&-3t \\
 sint&cost&0 \\
\end{vmatrix}=3ticost-3tjsint-(2cost+tsint)k"


"\frac{d}{dt}(G\times F)=(3cost-3tsint)i-(3sint+3tcost)j-(sint+tcost-2sint)k="

"=3(cost-tsint)i-3(sint+tcost)j-(tcost-sint)k"